### v1

@@ Abstract
Vaccines were critical in reducing hospitalizations and mortality during the COVID-19 pandemic Suthare069317 ; gupta2021vaccinations ; schneider2021hospitalizations . Despite their wide availability in the United States, 62% of Americans chose not to be vaccinated during 2021 mathieu2021global . While online misinformation about COVID-19 is correlated to vaccine hesitancy loomba2021measuring ; pierri2022online , little prior work has explored a causal link between real-world exposure to antivaccine content and vaccine uptake Allen24Science . Here we present a compartmental epidemic model that includes vaccination, vaccine hesitancy, and exposure to antivaccine content. We fit the model to observational data to determine that a geographical pattern of exposure to online antivaccine content across US counties is responsible for a pattern of reduced vaccine uptake in the same counties. We find that exposure to antivaccine content on Twitter caused about 750,000 people to refuse vaccination between February and August 2021 in the US, resulting in at least 29,000 additional cases and 430 additional deaths. This work provides a methodology for linking online speech to offline epidemic outcomes. Our findings should inform social media moderation policy as well as public health interventions.
@@
COVID-19, vaccine hesitancy, misinformation, causal model
Several socio-economic factors have been linked to vaccine hesitancy using correlational studies yasminCOVID19VaccineHesitancy2021 , but finding evidence of drivers remains a challenge. One potential driver is exposure to anti-vaccine content on social media. A laboratory study demonstrated that exposure to COVID-19 misinformation decreased willingness to be vaccinated loomba2021measuring . Using social media data, higher rates of misinformation were found to precede increases in COVID-19 infections in countries during early 2020 Gallotti2020 , though this may have been due to a general wave of COVID-19 discussion in early 2020 yangthecopvid2021 . A temporal correlation was also found between increased production of online vaccine misinformation and higher vaccine hesitancy, as well as lower uptake, across US states and counties pierri2022online .
These studies focus on associations and do not establish a causal link between real-world exposure to antivaccine content and vaccine uptake. A recent study found evidence that online exposure to vaccine-skeptical content causes vaccine hesitancy Allen24Science . The authors extrapolated from experiments based on a set of headlines to the predicted impact of URL views on Facebook. Here we measure the extent to which changes in levels of exposure to antivaccine content on Twitter in US counties caused changes not only in vaccine hesitancy, but also vaccination rates and COVID-19 cases. (Our analysis predates the platform’s name change to X.) To estimate the causal link between exposure to antivaccine tweets, vaccine hesitancy, reduced vaccine uptake, and cases, we extend the SIR epidemic model with states that represent vaccinated and vaccine-hesitant people. This model fits empirical data about COVID cases, vaccinations, and exposure to antivaccine tweets better than simpler models that ignore vaccine hesitancy. A key parameter is the rate at which people exposed to antivaccine content become vaccine hesitant. Fitting the model to the data yields a positive value for this parameter, indicating that increased exposure increases vaccine hesitancy. As more people become vaccine hesitant, vaccine uptake rates decline, leading to increased infections and deaths.
@@
We analyze the records of cases and vaccinations in US counties between February and August 2021 (see Methods). To connect these data to antivaccine content exposure, we identify and geolocate antivaccine tweets. We use a text classifier (see Methods) to identify COVID-related tweets as “Antivax” or “Other.” About 10% of tweets in our data set can be geolocated to specific US counties, and about 8% are identified as containing antivaccine content. This yields a dataset of 26 million geolocated tweets, 2.2 million of which are identified as containing antivaccine content. Figure 1 shows that antivaccine content increased around July 4th, 2021 and was broadly distributed throughout the US with limited geographic clustering; some counties produced orders of magnitude more antivaccine content per capita than others. To measure county-level exposure to antivaccine content, we combine the number of antivaccine tweets produced in each county with a network that captures the spread of this content from one county to another via retweets (see Methods).
@@ 1 Antivaccine Tweets Increase Vaccine Hesitancy
To infer the impact of antivaccine tweets on vaccine hesitancy, we model the COVID epidemic with an SIR-like compartmental model that we call SIRVA (see Methods). In addition to the standard Infected ($I$) and Recovered ($R$) compartments, the model has compartments for vaccinated people ($V$) and divides the Susceptible group ($S$) into those who are willing ($S^{\prime}$) and unwilling ($A$) to be vaccinated, see Figure 2a. The key epidemic model parameters are the infection rate $\beta$, the recovery rate $\rho$, the vaccination rate $\nu$, and the rate $\gamma$ at which people become unwilling to be vaccinated. The latter can be written as $\gamma=\gamma_{e}E+\gamma_{p}$, where $\gamma_{e}$ is the rate at which people become vaccine-hesitant per unit of exposure to antivaccine content ($E$) and $\gamma_{p}$ is the rate of conversion to vaccine hesitancy due to other factors. Finally, we express the vaccine-hesitant population at time $t$ as $A=\alpha_{t}S$, where the vaccine hesitancy ratio $\alpha_{t}$ changes by rate $\gamma$ each day and its initial value is an additional parameter $\alpha_{0}$.
Figure 2: SIRVA model. (a) Compartmental model diagram (see Methods). Note that $A=\alpha S$ and $S^{\prime}=(1-\alpha)S$, where $\alpha$ is the vaccine hesitancy ratio, i.e., the fraction of susceptibles who are unwilling to be vaccinated at time $t$. $E$ is the magnitude of exposure to antivaccine tweets and $\gamma_{e}$ is the rate at which people become vaccine-hesitant due to this exposure. (b) Posterior distribution of $\gamma_{e}$, with 95% high-density interval. The maximum-likelihood value is $\gamma_{e}\approx 0.21$. (c) Posterior distributions of other global model parameters. Note that there are multiple curves for $\nu$ and $\beta$ because we used different values of these parameters during different time periods to account for changing infectivity and national vaccine availability; these are ordered from earliest to latest. (d) Vaccinated population (V), and population who are both susceptible and vaccine-hesitant (A).
We apply this model to each county and use Bayesian Markov Chain Monte Carlo (MCMC) to infer the posterior distributions of the parameters from the data (see Methods). We are primarily interested in the parameter $\gamma_{e}$, which quantifies the impact of exposure on vaccine hesitancy. Inspecting the posterior distributions of the parameters (see Figure 2b,c), we find that $\gamma_{e}$ is greater than zero ($p=0.0002$) with an approximate magnitude of 0.21 and a 95% credible interval between 0.16 and 0.26. This indicates that increases in exposure to antivaccine content predict future increases in vaccine hesitancy, and subsequent decreases in the vaccine uptake rate.
@@ 2 Antivaccine Tweets Prevent Vaccinations
The SIRVA model does not specify a direct relationship between exposure and vaccine uptake rates, however we can use causal graphical modeling pearlDoCalculusRevisited2012 to assess this relationship (see Methods). We define the Average Treatment Effect (ATE) as the change in vaccinations per exposure to antivaccine tweets, where exposure is expressed in units of antivaccine tweets per capita. We derive an expression for the ATE and find its magnitude to be $-757$ (see Methods), with a 95% credible interval between $-575$ and $-931$ and high confidence that it is strictly less than zero ($p=0.0002$). This result allows us to infer that the relationship is causal; we consider potential confounding factors in the Discussion. We also compare these results to a simpler linear model relating exposure to vaccine uptake, and find a similar (but stronger) negative relationship; our method improves over the simpler linear model by accounting for additional causal confounders (see Discussion).
Based on the ATE, we estimate that antivaccine tweets induced approximately 750,000 people to refuse COVID vaccinations nationwide (with a 95% credible interval of 572,000–926,000) during the period from February 6th to August 9th, 2021 in the United States (see Methods). This represents only a small fraction of the total number of Americans who were unwilling to be vaccinated and were susceptible to infection ($A$), which we estimate at approximately 113 million people in August 2021, up from 86 million in February (see Figure 2d and Methods).
We can use measurements of vaccine effectiveness to provide a lower bound for the number of COVID cases and deaths that may have resulted from this reduction in vaccination. We estimate that among the 750,000 people who remained unvaccinated as a result of antivaccine tweets, there were about 29,000 COVID cases and 430 COVID-attributable deaths during February-August 2021 that would have been prevented without this additional vaccine hesitancy (see Methods). This represents a lower bound on the total impact because there would have been secondary infections outside the vaccine-hesitant population. Additionally, there would have been more cases and deaths following August 2021, which are not counted here.
To test if accounting for vaccine hesitancy improves model accuracy, we compared the predictions of our SIRVA model against a simpler model that does not include vaccine hesitancy (called SIRV, see Methods) using leave-one-out cross-validation and a Bayesian model fit score (see Methods). The standard errors of the score give a natural scale for comparing the relative accuracy of the models. The SIRVA score is more than two standard errors better than that of SIRV ($-(1698\pm 7)\times 10^{2}$ vs. $-(1716\pm 7)\times 10^{2}$, closer to zero is better). This indicates that SIRVA is more accurate at predicting unobserved data points.
@@ 3 Discussion
The proposed SIRVA model provides us with a novel approach to capture the role of vaccine hesitancy in epidemics. We use this model to measure the effect of antivaccine Twitter content on vaccine uptake during the COVID-19 pandemic in the United States. We find that exposure to antivaccine content on Twitter caused decreased vaccine uptake rates and increased cases and deaths.
@@
Our causal analysis hinges on a few key points. First and most critically, by leveraging the retweet network of COVID-related tweets between US counties, our measure of exposure to antivaccine content is specific to both the Twitter platform and the particular geographic distribution of Twitter users discussing COVID, allowing us to rule out many possible confounding factors that act through other social networks (e.g., Facebook) or with different geographic distributions. We test this geographic and platform specificity of our measure of exposure in two ways: (i) we tested the correlation of the COVID-related retweet network with other social networks (i.e., Meta’s Social Connectedness Index social_connectedness_index_Bailey2018 ), and found no correlation; and (ii) we tested whether shuffling exposure data by county would destroy the measured relationship between antivaccine tweets and increased vaccine hesitancy, and found that the relationship was null in this case.
@@
In addition to this platform and geographic specificity, we also accounted for other possible causal confounders in our model, including preexisting vaccine hesitancy in each county, nationwide drift in vaccine hesitancy overtime, and possible differential antivaccine content exposure rates based on preexisting vaccine hesitancy. We accounted for pre-existing antivaccine sentiments in each county by including an inferred free parameter for the initial antivaccine hesitancy ratio $\alpha_{0}$ in each county. We accounted for nationwide drift in vaccine hesitancy over time through an additional parameter $\gamma_{p}$ allowing for conversion to vaccine hesitancy without exposure to antivaccine tweets. Additionally, in our causal graphical model we accounted for the tendency of people with higher vaccine hesitancy to be more likely to be exposed to antivaccine content on social media by including an additional causal path from vaccine hesitancy to exposure, and this is reflected in our average treatment effect estimation. Finally, we also tested an alternative SIRVA model in which existing vaccine hesitancy may produce additional vaccine hesitancy within a county, for instance by word-of-mouth spread of antivaccine sentiment (see Methods); this more complex model produced qualitatively similar results as the SIRVA model, ruling out the possibility that word-of-mouth spread can explain the relationships we find.
@@
The accuracy of our method may be affected by data limitations. First, the official CDC data on COVID vaccinations, cases, and deaths contain imputed values and reporting lags. Second, our Twitter user geolocation data has limited coverage of the user population. Third, our observations may not capture the full spectrum of all antivaccine content on Twitter. These limitations may have created sample biases in the measured antivaccine content exposure and vaccine uptake rates.
With these caveats, our analysis estimates that 750,000 people became vaccine hesitant as a result of antivaccine content on Twitter. These are only a small fraction of our model’s estimate of 27 million Americans who became vaccine hesitant between February and August 2021. This larger increase is likely due to other sources of antivaccine messaging outside Twitter, including Facebook Allen24Science , traditional media, and word-of-mouth interactions. Further work could analyze data from other platforms, such as Instagram and TikTok, which also carry antivaccination content ortiz-sanchez_analysis_2020 .
This work constitutes a significant contribution to both public health research and social media studies because it establishes a causal link between online content and offline public health outcomes. These conclusions should inform future social media policy and epidemic modeling efforts.

### v2

@@ Abstract
Despite the wide availability of COVID-19 vaccines in the United States and their effectiveness in reducing hospitalizations and mortality during the pandemic, a majority of Americans chose not to be vaccinated during 2021. Recent work shows that vaccine misinformation affects intentions in controlled settings, but does not link it to real-world vaccination rates. Here, we present observational evidence of a causal relationship between exposure to antivaccine content and vaccination rates, and estimate the size of this effect. We present a compartmental epidemic model that includes vaccination, vaccine hesitancy, and exposure to antivaccine content. We fit the model to data to determine that a geographical pattern of exposure to online antivaccine content across US counties explains reduced vaccine uptake in the same counties. We find observational evidence that exposure to antivaccine content on Twitter caused about 14,000 people to refuse vaccination between February and August 2021 in the US, resulting in at least 510 additional cases and 8 additional deaths. This work provides a methodology for linking online speech with offline epidemic outcomes. Our findings should inform social media moderation policy as well as public health interventions.
@@
Vaccine, Misinformation, Causation, Hesitancy, COVID-19
1 Introduction
Vaccines were critical in reducing hospitalizations and mortality during the COVID-19 pandemic [1, 2, 3]. Despite their wide availability in the United States, 62% of Americans chose not to be vaccinated during 2021 [4]. Several socio-economic factors have been linked to vaccine hesitancy using correlational studies [5], but finding evidence of drivers remains a challenge. One potential driver is exposure to anti-vaccine content on social media [6]. A laboratory study demonstrated that exposure to COVID-19 misinformation decreased willingness to be vaccinated [7]. Using social media data, higher rates of misinformation were found to precede increases in COVID-19 infections in countries during early 2020 [8, 9], though this may have been due to a general wave of COVID-19 discussion in early 2020 [10]. A temporal correlation was also found between increased production of online vaccine misinformation and higher vaccine hesitancy, as well as lower uptake, across US states and counties [11]. A recent study combined experimental measurements of headlines’ effects on vaccination intentions with data on URL views to explore the causal link between exposure to vaccine-skeptical content on Facebook and vaccine hesitancy [12]. Many factors influence disease perceptions and vaccination decisions [13], so it is unclear how intention to vaccinate after exposure to online content translates into actual vaccine uptake.
To establish a causal link between real-world exposure to antivaccine content and vaccine uptake rates, we must begin by measuring these two specific variables. Here we measure the extent to which changes in levels of exposure to antivaccine content on Twitter (now X) in US counties caused changes not only in vaccine hesitancy, but also COVID-19 vaccination rates and cases. To estimate the causal link between exposure to antivaccine tweets, vaccine hesitancy, reduced vaccine uptake, and cases, we develop an original extension to the SIR epidemic model with states that represent vaccinated and vaccine-hesitant people. This model fits empirical data about COVID cases, vaccinations, and exposure to antivaccine tweets better than simpler models that ignore vaccine hesitancy. A key parameter is the rate at which people exposed to antivaccine content become vaccine hesitant. Fitting the model to the data yields a positive value for this parameter, indicating that increased exposure decreases vaccine uptake rates.
We analyze the records of cases and vaccinations in US counties between February and August 2021 (see Methods). To connect these data to antivaccine content exposure, we identify geolocated antivaccine tweets. We use a text classifier (see Methods) to identify COVID-related tweets as “Antivax” or “Other.” About 10% of all the tweets in our data set can be geolocated to specific US counties, and about 8% of the geolocated tweets are identified as containing antivaccine content. This yields a dataset of 26 million geolocated tweets, 2.2 million of which are identified as containing antivaccine content. Figure 1 shows that antivaccine content increased around July 4th, 2021 and was broadly distributed throughout the US with limited geographic clustering; some counties produced orders of magnitude more antivaccine content per capita than others. To measure county-level exposure to antivaccine content, we combine the number of antivaccine tweets produced in each county with a network that captures the spread of this content from one county to another via retweets (see Methods).
@@ 2 Antivaccine Tweets Increase Vaccine Hesitancy
To infer the impact of antivaccine tweets on vaccine hesitancy, we model the COVID epidemic with an SIR-like compartmental model that we call SIRVA (see Methods). In addition to the standard Infected ($I$) and Recovered ($R$) compartments, the model has compartments for vaccinated people ($V$) and divides the Susceptible group ($S$) into those who are willing ($S^{\prime}$) and unwilling ($A$) to be vaccinated, see Figure 2a. The key epidemic model parameters are the infection rate $\beta$, the recovery rate $\rho$, the vaccination rate $\nu$, and the rate $\gamma$ at which people become unwilling to be vaccinated. The latter can be written as $\gamma=\gamma_{e}E+\gamma_{p}$, where $\gamma_{e}$ is the rate at which people become vaccine-hesitant per unit of exposure to antivaccine content ($E$, see Methods) and $\gamma_{p}$ is the rate of conversion to vaccine hesitancy due to other factors. Finally, we express the vaccine-hesitant population at time $t$ as $A=\alpha_{t}S$, where the vaccine hesitancy ratio $\alpha_{t}$ changes by rate $\gamma$ each day and its initial value is an additional parameter $\alpha_{0}$.
Figure 2: SIRVA model. (a) Compartmental model diagram (see Methods). Note that $A=\alpha S$ and $S^{\prime}=(1-\alpha)S$, where $\alpha$ is the vaccine hesitancy ratio, i.e., the fraction of susceptibles who are unwilling to be vaccinated at time $t$. $E$ is the magnitude of exposure to antivaccine tweets and $\gamma_{e}$ is the rate at which people become vaccine-hesitant due to this exposure. (b) Posterior distribution of $\gamma_{e}$, with 95% high-density interval. The posterior mean value is $\gamma_{e}\approx 0.18$. (c) Posterior distributions of other global model parameters. Note that there are multiple curves for $\nu$ and $\beta$ because we used different values of these parameters during different time periods to account for changing infectivity and national vaccine availability (24-day periods for $\beta$, 48-day periods for $\nu$); these are ordered from earliest to latest. (d) Vaccinated population (V), and population who are both susceptible and vaccine-hesitant (A).
We apply this model to each county and use Bayesian Markov Chain Monte Carlo (MCMC) to infer the posterior distributions of the parameters from the data (see Methods). We are primarily interested in the parameter $\gamma_{e}$, which quantifies the impact of exposure on vaccine hesitancy. Inspecting the posterior distributions of the parameters (see Figure 2b,c), we find that $\gamma_{e}$ is greater than zero ($p=0.0002$) with an approximate magnitude of 0.18 and a 95% credible interval between 0.15 and 0.22. This indicates that increases in exposure to antivaccine content predict future increases in vaccine hesitancy, and subsequent decreases in the vaccine uptake rate.
@@ 3 Antivaccine Tweets Prevent Vaccinations
The SIRVA model does not specify a direct relationship between exposure and vaccine uptake rates, however we can use causal graphical modeling [14] to assess this relationship (see Methods).
We define the Average Treatment Effect (ATE) as the change in vaccinations per capita per unit exposure to antivaccine tweets, where exposure is expressed in units of antivaccine tweets per capita. We derive an expression for the ATE and find its magnitude to be $3.2\times 10^{-4}$ (see Methods). Note that the confidence that the ATE is less than zero is the same as the confidence that $\gamma_{e}$ is greater than zero ($p=0.0002$). Based on this ATE, we estimate that antivaccine tweets induced 14,086 people to refuse COVID vaccinations nationwide (with a 95% credible interval of 11,414 – 16,759) during the period from February 6th to August 9th, 2021 in the United States (see Methods). This represents only a small fraction of the total number of Americans who were unwilling to be vaccinated and were susceptible to infection ($A$), which we estimate at approximately 117 million people in August 2021, up from 98 million in February (see Figure 2d and Methods).
We can use measurements of vaccine effectiveness to provide a lower bound for the number of COVID cases and deaths that may have resulted from this reduction in vaccination. We estimate that among the 14,086 people who remained unvaccinated as a result of antivaccine tweets, there were about 511 COVID cases and 8 COVID-attributable deaths during February-August 2021 that would have been prevented without this additional vaccine hesitancy (see Methods). This represents a lower bound on the total impact because there would have been secondary infections outside the vaccine-hesitant population. Additionally, there would have been more cases and deaths following August 2021, which are not counted here.
To test if accounting for vaccine hesitancy improves model accuracy, we compared the predictions of our SIRVA model against a simpler model that does not include vaccine hesitancy (called SIRV, see Methods) using leave-one-out cross-validation and a Bayesian model fit score (see Methods). The standard errors of the score give a natural scale for comparing the relative accuracy of the models. The SIRVA model’s fit score is three standard errors better than that of SIRV ($(1716\pm 7)\times 10^{2}$ vs. $(1737\pm 7)\times 10^{2}$, closer to zero is better). This indicates that SIRVA is more accurate at predicting unobserved data points.
Coupled with our consideration of potential confounding factors (see Discussion), these results allow us to infer that the observed relationship between exposure antivaccine tweets and vaccine refusal is causal. We also compare our results to a simpler linear model relating exposure to vaccine uptake, and find a similar (but stronger) negative relationship; our method improves over the simpler linear model by accounting for additional causal confounders (see Discussion).
@@ 4 Discussion
The proposed SIRVA model provides us with a novel approach to capture the role of vaccine hesitancy in epidemics. We use this model to measure the effect of antivaccine Twitter content on vaccine uptake during the COVID-19 pandemic in the United States. We find evidence that exposure to antivaccine content on Twitter caused decreased vaccine uptake rates and increased cases and deaths. While this analysis was based on COVID, it may have implications for outbreaks of other vaccine-preventable diseases, such as measles.
@@
Our causal analysis hinges on a few key points. First and most critically, by leveraging the retweet network of COVID-related tweets between US counties, our measure of exposure to antivaccine content is specific to both the Twitter platform and the particular geographic distribution of Twitter users discussing COVID, allowing us to rule out many possible confounding factors that act through other social networks (e.g., Facebook) or with different geographic distributions. We test this geographic and platform specificity of our measure of exposure in two ways: (i) we tested the correlation of the COVID-related retweet network with other social networks (i.e., Meta’s Social Connectedness Index [15]), and found no correlation; and (ii) we tested whether shuffling exposure data by county would destroy the measured relationship between antivaccine tweets and increased vaccine hesitancy, and found that the relationship was null in this case.
@@
In addition to this platform and geographic specificity, we accounted for other possible causal confounders in our model, including preexisting vaccine hesitancy in each county, nationwide drift in vaccine hesitancy over time, and possible differential antivaccine content exposure rates based on preexisting vaccine hesitancy. We accounted for pre-existing antivaccine sentiments in each county by including an inferred free parameter for the initial antivaccine hesitancy ratio $\alpha_{0}$ in each county; this accounts for the impact that demographic and pre-existing political factors may have on antivaccine sentiment. We accounted for nationwide drift in vaccine hesitancy over time through an additional parameter $\gamma_{p}$ allowing for conversion to vaccine hesitancy without exposure to antivaccine tweets. Additionally, in our causal graphical model, we accounted for the tendency of people with higher vaccine hesitancy to be more likely to be exposed to antivaccine content on social media. To this end, we included an additional causal path from vaccine hesitancy to exposure, and this is reflected in our average treatment effect estimation. This explicitly accounts for any possible “reverse causation” feedback from vaccine hesitancy to exposure.
We also tested an alternative SIRVA model in which existing vaccine hesitancy may produce additional vaccine hesitancy within a county, for instance by word-of-mouth spread of antivaccine sentiment (see Methods); this more complex model produced qualitatively similar results as the SIRVA model, ruling out the possibility that word-of-mouth spread can explain the relationships we find. Between testing this alternative model and including the feedback from vaccine hesitancy to exposure in our causal graph, we believe we have accounted for any endogenous dynamics that could confound our measurement.
@@
There are different ways to model the effects of antivaccine content exposure on vaccination uptake. For instance, we could alternatively make the daily vaccination rate ($\nu$) dependent on exposure ($E$). This would be simpler than the SIRVA model, by eliminating the compartment A and the latent variables that are used to construct it. However, the SIRVA model lets us capture the intuition that vaccine hesitancy is a durable result of accumulated exposures over time rather than an immediate and transient effect of the most recent media exposures.
The accuracy of our method may be affected by data limitations and modeling assumptions. First, the official CDC data on COVID vaccinations, cases, and deaths contain imputed values and reporting lags. Second, our Twitter user geolocation data has limited coverage of the user population. Third, our observations may not capture the full spectrum of all antivaccine content on Twitter. These limitations may have created sample biases in the measured antivaccine content exposure and vaccine uptake rates.
On the modeling side, we neglect potential infections among vaccinated people, vaccinations among vaccine-hesitant people, and the lag time between exposure to COVID and active infection. In addition, while variations of epidemic models have been proposed to capture the spread of misinformation (e.g., [16, 17, 18]), we simply assume that the flow of antivaccine content between counties is proportional to retweet rates. Finally, we chose to measure an average treatment effect over all counties together, using a single key parameter $\gamma_{e}$ to capture the effect size, rather than attempting to measure the effect in each county individually or in smaller groupings of counties; this averaging was necessary to resolve a clear signal in our measurement, but makes the effect size seem homogeneous across all counties, when in fact it likely differs between counties. Future work could explore the heterogeneity of $\gamma_{e}$ across different (subsets of) counties. For example, one might expect different effects ($\gamma_{e}$ values) based on the political leanings of counties [11].
With these caveats, our analysis estimates that 14,086 people became vaccine hesitant as a result of antivaccine content on Twitter. These are only a very small fraction of our model’s estimate of 27 million Americans who became vaccine hesitant between February and August 2021. This larger increase in hesitancy is likely due to other sources of antivaccine messaging outside Twitter, including Facebook [12], traditional media, and word-of-mouth interactions. Further work could analyze data from other platforms, such as Instagram and TikTok, which also carry antivaccination content [19].
The impact of antivaccine content on public health outcomes highlighted by our analysis suggests that policy countermeasures may be necessary. Prior literature suggests a few possible approaches. Public health information campaigns may be used to “inoculate” the public against misinformation by spreading true information first [20, 21, 18], or to actively fact check misinformation that is already spreading [22, 23, 17]. Limiting the spread of misinformation on social media by removing posts or reducing their visibility can also be effective at reducing their impact [24, 17], especially if combined with other methods [25]. Similarly, removing “superspreaders” who propagate a large fraction of the misinformation on a platform can be an effective intervention [26]. In general, earlier interventions seem to work better than later ones [27].
This work contributes to both public health research and social media studies by establishing a causal link between online content and offline public health outcomes. These conclusions should inform future social media policy and epidemic modeling efforts.